Solve the following initial value problem 22 y (z) – 7 (de v(z)) + 10 y (2) = 18 sin (z) – 14 cos (2) with d. 2 y(0)=2, dy(0)could someone help me with this
Answer #1

consider the given IVP y-74tloy = 18 sina_14 asa with yo)=2 and yco)=12 28 5x the characteristic equation is mm +10=0 » m=9 A-7B = -14 and 7A+ 9B = 18 solving these equations, we get Azo and B=2 The particular solution is up=asina The general solu